OpenStudy (anonymous):

Find the derivative of f(x)=5/x at x = -1.

OpenStudy (anonymous):

by hand i bet

OpenStudy (anonymous):

I worked it out to [(5/x+h)-(5/x)]/h but I'm not sure if the 5 and x cancel out or not.

OpenStudy (anonymous):

\[f(x)=\frac5x\] Then : \[f'(x)=-\frac{5}{x^2}\] So : \[f'(-1)=-\frac{5}{(-1)^2}=-5\]

OpenStudy (anonymous):

Shouldn't you use the derivative equation

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

\[\lim_{x\to -1}\frac{\frac{5}{x}+5}{x+1} \]

OpenStudy (anonymous):

I don't know how to do that. What did you do there?

OpenStudy (anonymous):

if you are doing this by hand, not by shortcut methods, then \[\f'(a)=\lim_{x\to a}\frac{f(x)-f(a)}{x-a}\]

OpenStudy (anonymous):

since \(f(-1)=-5\) in your case you get \[f'(-1)=\lim_{x\to -1}\frac{\frac{5}{x}+5}{x+1}\]

OpenStudy (anonymous):

then it is algebra from here on in simplify the complex fraction to get \[\frac{\frac{5+5x}{x}}{x+1}=\frac{\frac{5(x+1)}{x}}{x+1}=\frac{5}{x}\]

OpenStudy (anonymous):

take the limit by replacing \(x\) by \(-1\) and get \(-5\) of course the shortcut method is quicker

OpenStudy (anonymous):

if you want to compute \[\frac{\frac{5}{x+h}-\frac{5}{x}}{h}\] we can do that too

OpenStudy (anonymous):

I'm doing precal so I haven't been introduced to any shortcuts. I think I understand now. Thank you for your help. I just thought you were supposed to use the derivative equation which is what I did. Can we try it that way

OpenStudy (anonymous):

I got to that part, but I don't know what to do from there

OpenStudy (anonymous):

it is just more algebra the numerator is \[\frac{5x-5(x+h)}{x(x+h)}=\frac{-5h}{x(x+h)}\]

OpenStudy (anonymous):

then when you divide by \(h\) the \(h\) cancels and you are left with \[\frac{-5}{x(x+h)}\]

OpenStudy (anonymous):

What happened to the second five?

OpenStudy (anonymous):

take the limit as \(h\to 0\) and get \(-\frac{5}{x^2}\) then replace \(x\) by \(-1\) and get \(-5\)

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

this is what you have to compute \[\lim_{h\to 0}\frac{\frac{5}{x+h}-\frac{5}{x}}{h}\]

OpenStudy (anonymous):

now lets ignore the \(h\) in the denominator for a moment, and concentrate only on \[\frac{5}{x+h}-\frac{5}{x}\]

OpenStudy (anonymous):

\[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\] so \[\frac{5}{x+h}-\frac{5}{x}=\frac{5x-5(x+h)}{x(x+h)}\] so far so good?

OpenStudy (anonymous):

So you cross multiplied and then multiplied the denominators?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just like with numbers right? \[\frac{5}{7}-\frac{3}{8}=\frac{5\times 8-7\times 3}{7\times 8}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we can compute \[\frac{5x-5(x+h)}{x(x+h)}=\frac{5x-5x-5h}{x(x+h)}=\frac{-5h}{x(x+h)}\]

OpenStudy (anonymous):

ok to there?

OpenStudy (anonymous):

okay. Now I see where the second five went. It got cancelled out. I'm ok

OpenStudy (anonymous):

now don't forget there was an \(h\) in the denominator of the original expression when you divide \(\frac{-5h}{x(x+h)}\) by \(h\) you are left only with \(-\frac{5}{x(x+h)}\)

OpenStudy (anonymous):

I got that. what next

OpenStudy (anonymous):

@satellite73

zepdrix (zepdrix):

So where'd you leave off? You were able to get a common denominator and combine those two fractions? \[\large \lim_{h \rightarrow 0} \frac{-5}{x(x+h)}\] Need help from this step?

OpenStudy (anonymous):

Yes!

zepdrix (zepdrix):

We can from this point plug zero directly in for h, since it no longer causes a problem for us.

OpenStudy (anonymous):

and then put -1 in for x^2?

zepdrix (zepdrix):

So we started with \(\large f(x)=\dfrac{5}{x}\) By using the limit definition of the derivative we were able to determine that \(\large f'(x)=-\dfrac{5}{x^2}\) And now they're asking for \(\large f'(-1)\). Yes very good :) plug away!

OpenStudy (anonymous):

Thank you so much! Would you like to help me with another problem? It involves finding a limit algebraically.

zepdrix (zepdrix):

Close this thread, it's gotten rather long and messy. Open a new one with your new question. Yah I'll come take a look :)

OpenStudy (anonymous):

thank you!